OpenStudy (anonymous):

One number is 4 more than three times a smaller number. If twice the larger number is decreaased by three times the smaller number, the result is 32. Find the number. I know the answer but i need help solving it because i have a test tomorrow thanks!

OpenStudy (inkyvoyd):

Okay. One number is 4 more than 3 times a smaller number. Sounds like we have two numbers. So let's set x=larger number y=smaller number.

OpenStudy (anonymous):

i did x= the smaller 2x+4= the larger

OpenStudy (inkyvoyd):

Okay. Then x is the smaller, and 2x+4 is the larger.

OpenStudy (inkyvoyd):

Are you learning systems of equations? or just one variable?

OpenStudy (anonymous):

we learned system of equations last year

OpenStudy (inkyvoyd):

Okay. I'll straight up tell you that while it might be possible to solve this with just x as a variable, it will be much less confusing to do with x and y.

OpenStudy (anonymous):

okay so ill use x and y thankss

OpenStudy (inkyvoyd):

Alright. Now really the only reason I set x as the larger number is because it turned up first. y as the smaller number cause it turned up second.

OpenStudy (inkyvoyd):

We are looking for two equations that relate x and y. can you figure out the first?

OpenStudy (anonymous):

3x+4?

OpenStudy (inkyvoyd):

mm. Close, but remember that x is the larger number. You should be getting 3y+4=x

OpenStudy (anonymous):

wait no 2(3x+4)-3x=32

OpenStudy (anonymous):

ohh wait what?

OpenStudy (inkyvoyd):

One number is 4 more than three times a smaller number.

OpenStudy (inkyvoyd):

We have a number x that is 4 more than 3 times a smaller number.

OpenStudy (anonymous):

ok i got that but then whats the second?

OpenStudy (inkyvoyd):

and the smaller number is y. so we have x is 4 more than 3 times y. or, x=4+3y in other words 3y+4=x

OpenStudy (inkyvoyd):

If twice the larger number is decreaased by three times the smaller number, the result is 32.

OpenStudy (inkyvoyd):

Replace the larger number and smaller number to what we have set them to, x and y ;) Tell me what you get...

OpenStudy (anonymous):

2x-3y=32

OpenStudy (inkyvoyd):

Yes. Good! Now we set up the system. 3y+4=x 2x-3y=32

OpenStudy (anonymous):

substitution

OpenStudy (inkyvoyd):

yup.

OpenStudy (anonymous):

i got 10.6 repeting

OpenStudy (inkyvoyd):

Can you show your work? that's not the right answer :S

OpenStudy (anonymous):

@inkyvoyd 2(3y+4)-3y=32 6y+8-3y=32 3y+8=32 -8 -8 -------------- 3y=32 --- --- 3 3 y=10.66666666666666666666666666

OpenStudy (inkyvoyd):

3y+8=32 catch: 3y+8=24

OpenStudy (inkyvoyd):

y=3

OpenStudy (anonymous):

so x is 13:)

OpenStudy (inkyvoyd):

No, check again :)

OpenStudy (anonymous):

3(3)+4=x 13=x

OpenStudy (anonymous):

oh i get itt

OpenStudy (anonymous):

thankss sooo muchh

OpenStudy (inkyvoyd):

yeah sorry, I made a mistake to. y=8 lol

OpenStudy (anonymous):

TANK YOU

OpenStudy (inkyvoyd):

No problem :)